# Density Descent for Diversity Optimization

### v1

@@ Appendix F Statistical Analysis
Table 2: One-way ANOVA results in each domain. All $p$-values are less than 0.001.
@@
Table 3: Pairwise comparisons for coverage in each domain. Each entry compares the method in the row to the method in the column. For instance, we can see that DDS-KDE had significantly higher coverage than NS in LP. Comparisons were performed with Tukey’s HSD test. $>$ indicates significantly greater, $<$ indicates significantly less, and -- indicates no significant difference. $\varnothing$ indicates an invalid comparison.
@@
Table 4: Pairwise comparisons for cross-entropy in each domain. Note that lower cross-entropy is better, so significantly less ($<$) indicates that a method is significantly better.
@@
Appendix G Heatmaps
Figure 5: Heatmaps of DDS-KDE, DDS-CNF, NS, and CMA-MAE in the three domains that have 2D feature spaces (LP, Arm Repertoire, Deceptive Maze).
i2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ Appendix F Statistical Analysis
Table 2. One-way ANOVA results in each domain. All $p$-values are less than 0.001.
@@
Table 3. Pairwise comparisons for coverage in each domain. Each entry compares the method in the row to the method in the column. For instance, we can see that DDS-KDE had significantly higher coverage than NS in LP. Comparisons were performed with Tukey’s HSD test. $>$ indicates significantly greater, $<$ indicates significantly less, and -- indicates no significant difference. $\varnothing$ indicates an invalid comparison.
@@
Table 4. Pairwise comparisons for cross-entropy in each domain. Note that lower cross-entropy is better, so significantly less ($<$) indicates that a method is significantly better.
@@
Appendix G Coverage Map in Feature Space
Figure 5. Heatmaps of DDS-KDE, DDS-CNF, NS, and CMA-MAE in the three domains that have 2D feature spaces (LP, Arm Repertoire, Deceptive Maze).